# SeeThruFinger: See and Grasp Anything with a Multi-Modal Soft Touch

### v3

@@ 5.1 Markerless, Real-Time & Large-Scale Deformation Tracking of the See-Thru-Network
SeeThruFinger generically describes a class of soft robotic fingers/fingertips that involves a See-Thru-Network with in-finger vision as a design feature. The See-Through feature enables the in-finger vision to simultaneously capture visual features from the finger structure and the external environment. In this study, the See-Thru-Network is implemented using the Soft Polyhedral Networks. However, researchers have a wide range of design choices to customize the soft interface as needed, as long the resultant image contains sufficient details from both the finger structure and the external environment. The resultant performance is determined by the level of detail when extracting the soft structure’s deformation and the reusable information left to reconstruct the scene.
@@
Algorithmic development in video object segmentation is a well-researched field that enables convenient implementation of this capability for See-Thru-Networks with in-finger vision. The 30 Hz mask tracking implemented in this study is sufficiently high to enable real-time processing for robotics without adding to the computational burden of the robotic systems. Future research to allow a higher mask tracking bandwidth is preferred to increase the upper bound in collecting sensory feedback (towards 1,000 Hz) from in-finger vision, requiring higher framerate miniature cameras and high-performance mask tracking algorithms for more challenging robotic applications. The quality of mask segmentation at this step generally determines the maximum performance one can expect from the SeeThruFinger architecture in multi-modal perception. The solution to this problem provides the foundation to implement suitable robotic solutions for the second and third research problems involved in SeeThruFingers.
@@
5.2 Multi-Modal Visual Perception of the Occuluded Scene via In-Finger Vision
@@
An intuitive understanding of the second research problem is whether we can use in-finger vision for environmental perception, which is multimodal by nature. This eliminates the need for an external camera in modern robot workstation setups.
@@
Using in-finger vision for environmental perception is a counterintuitive application that can potentially remove the need for external cameras in a typical setup of robot workstations. Having an Eye camera fixed to the robot base is generally perceived as bio-inspired by humans or most living creatures, where visual organs are on the body for environmental perception, which is also shared by most robotic systems. After masked segmentation of the See-Thru-Network, the original image becomes fragmented. Still, it could be alternatively repaired using inpainting, a highly active field of research in computer vision and widely adopted practice throughout history. A direct result of inpainting the in-finger vision is the reconstruction of the scene, producing visual perceptual capabilities almost precisely as a typical Eye camera. It should be noted that the inpainting technique is computationally intensive, requiring high-performing hardware and long processing time, neither of which is ideal for robotic applications. In this study, we added an extra head-up motion during motion planning to collect an extra 1-second video clip for inpainting as a trade-off.
@@
In the sections above, we demonstrated object detection using inpainted images from in-finger vision. Here, we showcase the performance of achieving other modalities, such as depth sensing and scene segmentation, using the inpainted images, which are also widely used in modern robotics. In principle, robotists can reuse the inpainted images for any desirable extra modalities as needed in the same way as a computer vision problem, making the SeeThruFinger architecture a generic solution for robotics where constraints in hardware cost and computational power are of significant consideration.
5.2.1 Depth Sensing
@@
As a proof-of-concept, we conducted the monocular depth estimation based on the inpainted scene. We compared the reconstructed scene depths from both the in-finger camera and a camera with an unobstructed view of the robot’s working space. In this experiment, ten objects with various shapes and colors (milk powder can, tape, rubber ball, strawberry, glue, plum, glue stick, power drill with a packaging box, and drill toolkit) were placed on the tabletop. The robotic arm swept across the tabletop from left to right and returned to the starting point with a fixed orientation while the in-finger camera inspected the objects. We recorded both the in-finger and unobstructed videos by removing the SPN from the gripper. The former was further inpainted by E2FGVI algorithm. The inpainted and unobstructed videos are then fed to the Depth Anything V2 model [44] to obtain the depth estimations. The pixel-wise estimations are relative depth values, providing helpful information on the relative poses of the objects on the table. Hence, instead of comparing the actual values of the depth map, we adopt the structural similarity index (SSI) to evaluate the similarity between the depth map reconstructed from the inpainted view and that from the unobstructed view.
@@
5.2.2 Scene Segmentation
@@
Furthermore, we explored the feasibility of multi-object segmentation on the inpainted scene using the latest Segment Anything Model 2 (SAM 2) [45]. We tested on the same scene described in the depth estimation task. As shown in Fig. 13A, segmentation on the inpainted view achieves satisfactory performance with J measurement ranging between 0.64 and 0.87, which is 16% lower than the performance of segmentation on the unobstructed view on average. An interesting observation occurred in the segmentation of object 9 (a power drill with a packaging box), where the inpainted view performed better than the unobstructed view. Further frame-by-frame examination shows that the worse segmentation in the unobstructed view was caused by the intense surface reflection of the drill’s packaging box, and in some frames, SAM2 failed to segment object 9. On the contrary, SAM2 kept tracking object nine well in the inpainted view (see supplementary video). The snapshots of the unobstructed and in-finger vision and segmentations of the inpainted scene at two different times are shown in Figs. 13C&D.
5.3 Multi-Modal & Markerless Vision-based Tactile Sensing with the SeeThruFinger
@@
An intuitive understanding of the third research problem is whether we can use in-finger vision with the SeeThruFingers to achieve high-performing, multi-modal tactile sensing in a markerless way. Similar capabilities have been widely researched in the past literature, but the involvement of artificial markers remains a challenge among the current solutions [46, 32, 47, 48]. In this study, a by-product while addressing the first research problem is the real-time, pixel-wise segmentation of the See-Thru-Network’s mask throughout the implementation of the SeeThruFinger architecture. Visual features from this mask provide a robust sensory input to infer force-and-torque information for tactile grasping once a calibration dataset is obtained against commercial force-and-torque sensors for model training.
@@
While it is reasonable to use simpler neural networks, we adopted a generative approach in this study and utilized the Supervised Variational Autoencoder (SVAE) to address this problem. The latent variables obtained from SVAE are a powerful tool to infer various sensory modalities, requiring little disruption to the overall computational architecture. Besides the 6D force-and-torque, we demonstrate a simple implementation to infer the contact event on the SeeThruFinger. Recent research also shows promising results in proprioceptive shape reconstruction, contact point estimation, and tactile sensing using in-finger vision with the soft polyhedral networks used in this study, which could be further implemented using the deformation mask tracked in this study.
5.3.1 Contact Event Detection
@@
In previous work where the complex deformations of SPN are physically encoded into a much lower 6D space of AruCo Marker pose, the contact location information has been lost [33]. With a much higher dimension of 32 in the SVAE framework, We hypothesize such information could be maintained in the latent space to a certain degree.
A preliminary test was conducted by categorizing the training samples into three contact events: no contact, contact at the finger’s tip, and middle regions. We trained a contact event classifier using the SVAE framework shown in Fig. 9, which can be formulated by
@@
Figure 14 shows the confusion matrix of the classifier evaluated on the validation dataset of 8000 samples. The accuracies for all three contact events are above 97%. Considering the classifier only takes the SPN’s mask as input, which filters out all the see-through visual information, the results confirm our hypothesis that the soft touch represented by the latent vectors contains tactile details such as resultant force/torque and contact location. A real-time demonstration is included in the video supplementary. The binary classes of the contact location are partly limited by the data collection process, where the precise locations were not recorded. The granularity might be further improved by having more dense labels or even becoming a regression task.
5.4 Multi-Modal Perceptual Embodiment via a Soft Touch
Using a single sensory input from the in-finger vision of a SeeThruFinger featuring See-Thru-Networks, this study presents a multi-modal approach towards robotic perceptual embodiment via vision-based tactile sensing with a soft touch, as shown in Fig. 15. Our proposed SeeThruFinger architecture generates various outputs for extrinsic perception, including inpainted scenes, object detection, depth sensing, and scene segmentation. For intrinsic perceptions, we have demonstrated the implementation of the SeeThruFinger architecture that generates outputs including masked deformation tracking, 6D force and torque sensing in real-time and high accuracy, and contact event detection. Further expansion to other intrinsic or extrinsic modalities can be implemented depending on the specific need in robotics or applications.
6 Conclusion & Future Works
This paper presents the SeeThruFinger architecture for Multi-Modal and Vision-Based Tactile Sensing via a See-Thru-Network, simultaneously achieving multi-modal visual perception and tactile sensing through a single in-finger vision of the Soft Polyhedral Networks. The occluded in-finger vision provides sufficient scene details that could be algorithmically inpainted for visual perception. This is achieved by tracking a deformable mask of the Soft Polyhedral Network during contact-based interactions using real-time video object segmentation. The omni-directional adaptation of the Soft Polyhedral Network provides a rich set of visual features that could be directly used to learn a tactile sensing model for reactive grasping. Using masked Soft Polyhedral Networks as the visual input effectively removes the necessity for artificial markers in vision-based tactile sensing. The resultant SeeThruFinger architecture enables multi-modal perception of the scene, including globalized scene segmentation object detection and localized tactile sensing at high accuracy. As reported in this study, a direct benefit is the possibility of completely removing the external camera while simultaneously providing physical compliance, visual perception, and tactile sensing using just one in-finger vision. Our proposed SeeThruFinger architecture should also apply to other vision-based tactile sensing solutions with an illuminated chamber by removing the opaque coating on the outer surface of the soft membrane [28]. However, the performance may be reduced due to the blurry image of the scene occluded by the transparent soft membrane, which may only work at a very close range that is insufficient for visual perception of the whole scene [49].
Our work has several limitations. Even though the XMem is reasonably fast (30 Hz), it is still far from the chosen camera’s 330 Hz framerate. The whole pipeline is not as efficient yet, which requires further optimization for robotic interactions. In this work, we did not fine-tune the object detection model for the objects tested, and the object detection results show mislabeled object classes even though they can be detected. The learned tactile perception model is subject to a systematic shift when applied to a new finger that is not presented in the training dataset due to the slight difference in the initial template mask of each finger. Although we fed the initial template mask along with the instantaneous one, SVAE failed to encode the geometric deformation compared to the template mask, probably because it only saw two templates during training. Besides mitigating this systematic error by simple subtraction, a promising way is through data augmentation.
There are a few directions we would like to explore further in the future. One is underwater visual perception and tactile sensing using SeeThruFinger by translating models learned on land to underwater scenarios, aiming at dexterous and sensitive exploration underwater [50]. Due to the unique design, the SeeThruFinger can be directly used underwater if the camera housing is waterproofed [36], which is quite simple as there are no dynamic seals. Another potential application is embodied robot learning with multi-modal tactile perception, where the SeeThruFingers can be a promising solution to introduce simultaneous benefits in omni-directional adaptation in grasping while providing a rich set of perceptual modalities including high-performing 6D force-and-torque sensing, which are yet to be implemented with ease for systems such as the Universal Manipulation Interface [51]. In this study, only one of the SeeThruFingers’ in-finger vision was used, but both are fully functional. It would be interesting to see if the stereo vision for 3D reconstruction or 6D pose estimation based on the inpainted scene from both in-finger visions would be possible. This is another exciting direction we would like to explore further.
@@ Acknowledgement
This work was partly supported by the National Natural Science Foundation of China under Grant 62206119 and Grant 62473189, the Science, Technology, and Innovation Commission of Shenzhen Municipality under Grant JCYJ20220818100417038, Shenzhen Long-Term Support for Higher Education at SUSTech under Grant 20231115141649002 and SUSTech Virtual Teaching Lab for Machine Intelligence Design and Learning under Grant XJZLGC202241.
@@
Please refer to this link for a supplementary video of this paper.

### v4

@@ 4.2.1 Markerless, Real-Time & Large-Scale Deformation Tracking
VBSeeThruP generically describes a class of soft robotic fingers/fingertips that involves a See-Thru-Network with in-finger vision as a design feature. The See-Through feature enables the in-finger vision to simultaneously capture visual features from the finger structure and the external environment. In this study, the See-Thru-Network is implemented using the Soft Polyhedral Network. However, researchers have a wide range of design choices to customize the soft interface as needed, as long as the resultant image contains sufficient details from both the finger structure and the external environment. The resultant performance is determined by the level of detail when extracting the soft structure’s deformation and the reusable information left to reconstruct the scene.
@@
Algorithmic development in video object segmentation is a well-researched field that enables convenient implementation of this capability for See-Thru-Networks with in-finger vision. The 30 Hz mask tracking implemented in this study is sufficiently high to enable real-time processing for robotics without adding to the computational burden of the robotic systems. Future research to allow a higher mask tracking bandwidth is preferred to increase the upper bound in collecting sensory feedback (towards 1 kHz) from in-finger vision, requiring higher-framerate miniature cameras and high-performance mask tracking algorithms for more challenging robotic applications. The quality of mask segmentation at this step generally determines the maximum performance one can expect from the VBSeeThruP architecture in multi-modal perception. The solution to this problem provides the foundation for implementing suitable robotic solutions to address the second and third research problems.
@@
4.2.2 Multi-Modal Visual Perception of the Occluded Scene via In-Finger Vision
@@
An intuitive understanding of the second research problem is whether we can utilize in-finger vision for environmental perception, which is inherently multi-modal, thereby eliminating the need for an external camera in modern robot workstations.
@@
Using in-finger vision for environmental perception is counterintuitive. Having an Eye camera fixed to the robot base is generally perceived as bio-inspired by humans or most living creatures, where visual organs are on the body for environmental perception. This is also a feature shared by most robotic systems. After masked segmentation of the See-Thru-Network, the original image becomes fragmented. Still, it could be alternatively repaired using inpainting, a highly active field of research in computer vision and widely adopted practice throughout history. A direct result of inpainting the in-finger vision is the reconstruction of the scene, producing visual perceptual capabilities almost precisely as a typical Eye camera. It should be noted that the inpainting technique is computationally intensive, requiring high-performing hardware and long processing time, neither of which is ideal for robotic applications. In this study, we added an extra head-up motion during motion planning to collect an extra 1-second video clip for inpainting as a trade-off.
@@
In the sections above, we demonstrated object detection using inpainted images from in-finger vision. Here, we showcase the performance of achieving other modalities, such as depth sensing and scene segmentation, using the inpainted images, which are also widely used in modern robotics. In principle, robotists can reuse the inpainted images for any desirable extra modalities as needed, just as in a computer vision problem, making the VBSeeThruP architecture a generic solution for robotics, where constraints in hardware cost and computational power are significant considerations.
Depth Sensing
@@
We estimated monocular depth based on the inpainted scene as proof of concept. We compared the reconstructed scene depths from both the in-finger camera and a camera with an unobstructed view of the robot’s working space. In this experiment, ten objects with various shapes and colors (milk powder can, tape, rubber ball, strawberry, glue, plum, glue stick, power drill with a packaging box, and drill toolkit) were placed on the tabletop. The robotic arm swept across the tabletop from left to right and returned to the starting point with a fixed orientation while the in-finger camera inspected the objects. We recorded both the in-finger and unobstructed videos by removing the SPN from the gripper. The former was further inpainted by the E2FGVI algorithm. The inpainted and unobstructed videos are then fed to the Depth Anything V2 model [71] to obtain the depth estimations. The pixel-wise estimations provide relative depth values, offering helpful information on the relative poses of objects on the table. Hence, instead of comparing the actual values of the depth map, we adopt the structural similarity index (SSI) to evaluate the similarity between the depth map reconstructed from the inpainted view and that from the unobstructed view.
@@
Scene Segmentation
Furthermore, we explored the feasibility of multi-object segmentation on the inpainted scene using the latest Segment Anything Model 2 (SAM2) [72]. We tested on the same scene described in the depth estimation task.
@@
As shown in Fig. 12A, segmentation on the inpainted view achieves satisfactory performance with J measurement ranging between 0.64 and 0.87, which is 16% lower than the performance of segmentation on the unobstructed view on average. An interesting observation occurred in the segmentation of object 9 (a power drill with a packaging box), where the inpainted view performed better than the unobstructed view. Further frame-by-frame examination shows that the worst segmentation in the unobstructed view was caused by the intense surface reflection of the drill’s packaging box, and in some frames, SAM2 failed to segment object 9. On the contrary, SAM2 successfully tracked object nine in the inpainted view. The snapshots of the unobstructed and in-finger vision and segmentations of the inpainted scene at two different times are shown in Figs. 12C&D.
Multi-Modal & Markerless VBSeeThruP
@@
An intuitive understanding of the third research problem is whether we can utilize in-finger vision with a soft finger to achieve high-performing, multi-modal perception in a markerless manner. Similar capabilities have been widely researched in the past literature, but the involvement of artificial markers remains a challenge among the current solutions [73, 74, 75, 76]. In this study, a by-product while addressing the first research problem is the real-time, pixel-wise segmentation of the See-Thru-Network’s mask throughout the implementation of the VBSeeThruP architecture. Visual features from this mask provide a robust sensory input to infer FT information once a calibration dataset is obtained against commercial FT sensors for model training.
@@
While it is reasonable to use simpler neural networks, we adopted a generative approach in this study and utilized the Supervised Variational Autoencoder (SVAE) to address this problem. The latent variables obtained from SVAE are a powerful tool for inferring various sensory modalities, requiring minimal disruption to the overall computation. Besides the 6D FT, we demonstrate a simple implementation to infer the contact event on the soft finger. Recent research also shows promising results in proprioceptive shape reconstruction [17], contact point estimation [34], and tactile perception using in-finger vision [18] with the soft polyhedral network used in this study, which could be further implemented using the deformation mask tracked in this study.
Contact Event Detection
@@
In previous work where the complex deformations of SPN are physically encoded into a much lower 6D space of AruCo Marker pose, the contact location information has been lost [18]. With a much higher dimension of 32 in the SVAE framework, we hypothesize that such information could be maintained in the latent space to a certain degree. A preliminary test was conducted by categorizing the training samples into three contact events: no contact, contact at the finger’s tip, and the middle region. We trained a contact event classifier using the SVAE framework shown in Fig. 7, which can be formulated by
@@
Figure 13 shows the confusion matrix of the classifier evaluated on the validation dataset of 8,000 samples. The accuracies for all three contact events are above 97%. Considering the classifier only takes the SPN’s mask as input, which filters out all the see-through visual information, the results confirm our hypothesis that the soft touch represented by the latent vectors contains information such as force/torque and contact location. A real-time demonstration is included in the supplementary video. The binary classes of the contact location are partly limited by the data collection process, where the precise locations were not recorded. The granularity could be further improved by using denser labels or by converting it into a regression task.
5 Conclusion, Limitations, and Future Directions
Inspired by sensory substitution from biological agents, this study presents a Vision-based See-Through Perception (VBSeeThruP) approach to reason among different image layers for visual- and contact-based information from a single camera input. By placing a See-Thru-Network in front of a single in-finger camera, two layers of information can be obtained: one at a closer range for the soft interface, awaiting contact interactions, and another about the scene at a distance but occluded. Contact-based perception is substituted markerlessly by reasoning about the masked deformation, facilitating reactive see-through grasping.
Despite encouraging results, there are notable constraints. First, while XMem runs at 30 Hz, it remains significantly slower than the camera’s potential 330 Hz frame rate, and the overall pipeline still needs optimization for high-speed robotic tasks. Second, mislabeled detections sometimes occur because we did not fine-tune object detection for specific targets. Third, the learned 6D force/torque model exhibits a systematic shift when applied to new fingers with slightly different initial templates. Although subtracting this offset mitigates the error, more effective data augmentation strategies could further improve generalization. Finally, larger scenes often appear blurry through the STN, limiting the range of reliable visual perception.
Several directions remain open. Promising avenues include underwater perception and manipulation [77], embodied robot learning with multi-modal feedback that integrates SPN’s omni-directional adaptability and precise 6D force/torque sensing [78], and extending the single-camera approach to stereo vision for improved 3D reconstruction or 6D pose estimation.
@@ Acknowledgement
This work was partly supported by the National Natural Science Foundation of China under Grants 62473189 and 62206119, as well as the Shenzhen Long-Term Support for Higher Education at SUSTech under Grant 20231115141649002.
@@
Please refer to this link for supplementary videos of this paper.